You have to solve the equations using matrices. I can get part of the answer correct but not the whole answer. I can solve for y, but what I come up with for x is not correct according to the book. The two equations are:
12x + 2y = 2 and 6x - 3y = -11
I can come up with 3 for y but I end up with -11/6 for x and the book says it should be -1/3. I can't figure out how to get the -1/3.

Answer by Edwin McCravy(20063)

: You have to solve the equations using matrices. I can get part of the answer correct but not the whole answer. I can solve for y, but what I come up with for x is not correct according to the book. The two equations are:
12x + 2y = 2 and 6x - 3y = -11
I can come up with 3 for y but I end up with -11/6 for x and the book says it should be -1/3. I can't figure out how to get the -1/3.

12x + 2y =   2
 6x - 3y = -11

[12   2 |   2]
[ 6  -3 | -11]

Get a 0 where the 6 is by adding -1 
times the top row to 2 times the 
bottom row:

-1[12   2 |   2]
 2[ 6  -3 | -11]

  [12   2 |   2]
  [ 0  -8 | -24]

 
Get a 0 where the first 2 is in the top
row by adding 1 times the bottom
row to 4 times the top row

 4[12   2 |   2]
 1[ 0  -8 | -24]

  [48   0 | -16] 
  [ 0  -8 | -24]

Now divide the top row thru by 48
and the bottom row by -8:

  [1  0 | -1/3]
  [0  1 |   3 ]

So the solution is x = -1/3, y = 3

Edwin